OpenStudy (anonymous):

what is the opposite of the number 5 1/2???

OpenStudy (anonymous):

2/51

OpenStudy (sepeario):

Improper fraction = 11/2 flip it over = 2/11

OpenStudy (sepeario):

Called the 'recoprical'.